# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

### v2

@@ 5.3.2 Performance Under Distribution Shift
The distribution shift evaluations are provided in Table 2. UQ4CT provides substantial improvements in terms of ECE while maintains similar accuracy for both smaller and larger distribution shift scenarios. For smaller distribution shifts, UQ4CT shows comparable ECE performance as the in-distribution scenario. For the more challenging larger distribution shifts, UQ4CT still achieves the best ECE performance among all baseline models. Note that UQ4CT also achieves competitive prediction accuracy across all domain-specific tasks. This empirically shows that our proposed alignment of the functional epistemic uncertainty with predictive correctness improves generalizability and mitigating the overconfidence problem on the fine-tuned model.
@@
Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintaining similar accuracy.
@@
5.4 Ablation Study
In this section, we conduct ablation studies to investigate the effectiveness of our designed calibration loss, $\mathcal{L}_{cal}$. We first evaluate the incremental weighting performance of the calibration term, which investigates the effectiveness of $\mathcal{L}_{cal}$ at the early stage of fine-tuning. Then, we perform a sensitivity test, where we explore the overall performance impact of $\mathcal{L}_{cal}$. We also conduct an ablation study on the impact of active LoRA experts in Appendix A.4.
5.4.1 Incremental Weighting on Calibration Term
As shown in Table 3, the incremental loss has significantly worse ECE performance across all tasks. This demonstrates the advantage of uncertainty calibration even in the early stage. In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high epistemic uncertainty. Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
$\beta$ ACC $\uparrow$ ECE $\downarrow$
0 $75.5_{2.91}$ $12.9_{1.99}$
0.2 $76.0_{0.6}$ $7.47_{0.78}$
0.5 $75.9_{0.31}$ $7.82_{0.93}$
0.8 $76.6_{0.4}$ $5.94_{1.16}$
1 $\mathbf{77.3_{1.36}}$ $\mathbf{5.0_{1.15}}$
5.4.2 Sensitivity Test on Calibration Term
To further understand the effectiveness of the calibration loss, we perform a sensitivity test of the coefficient $\beta$ in Equation 14. This evaluates how our proposed calibration of parameter mixtures affect the overall model prediction and uncertainty quantification capabilities. We evaluate $\beta$ values of $0,0.2,0.5,0.8$ and $1$, where $\beta=0$ resembles the original MixLoRA method.
6 Discussion & Conclusion
In this work, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which addresses the overconfidence issues commonly encountered during fine-tuning of large language models. We present a functional perspective on quantifying epistemic uncertainty in LLMs and utilize it for uncertainty-calibrated fine-tuning. By incorporating functional-level epistemic uncertainty quantification with a mixture-of-experts framework, our proposed uncertainty-calibrated training loss effectively addresses the challenge of overconfidence in fine-tuned LLMs by significantly improving uncertainty calibration while maintaining high accuracy. Our evaluations demonstrate that UQ4CT reduces the Expected Calibration Error by more than 25% without compromising accuracy across a variety of downstream tasks, including common-sense and domain-specific reasoning, under in-distribution and out-of-distribution scenarios.
Acknowledgement
Appendix A Appendix
A.1 Load Balancing Loss
where
A.2 Experimental Results with Mistral-7B
@@
A.3 Incremental Weighting on Calibration Term
@@
As shown in Table 3, the incremental loss has significantly worse ECE performance across all tasks. This demonstrates the advantage of uncertainty calibration even in the early stage. In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high epistemic uncertainty. Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
A.4 Deciding Number of Active Experts
@@
A.5 Training Details
We train our model with total of $8$ LoRA experts, and select $2$ experts with the highest probability. For each expert, we use $rank=16$ and $alpha=32$. We use batch size of $16$ to train our model. For climate task, we set the learning rate to $45e-$ 4 and dropout rate to $0.1$ to incorporate the small dataset size. For other tasks, we use $42e-$ 4 as our learning rate with dropout $0.05$. We use AdamW as our optimizer and a cutoff length of $512$ for prompts during training.
@@ A.5 Training Details
A.6 Expected Calibration Error
@@
A.7 MMLU Distribution Shift Dataset Composition

### v3

@@ 5.2 Performance Under Distribution Shift
Table 1 presents the distribution shift evaluations. UQ4CT achieves substantial ECE improvements while maintaining high accuracy across both smaller and larger distribution shifts. For smaller shifts, UQ4CT’s ECE remains comparable to the in-distribution scenario. Under larger shifts, UQ4CT attains the lowest ECE among all baselines and delivers competitive accuracy on all domain-specific tasks. These results demonstrate that aligning uncertainty at the functional level with predictive correctness improves generalizability and mitigates overconfidence in fine-tuned models.
5.3 Ablation Studies
We conduct ablation studies to investigate the effectiveness of our designed calibration loss, $\mathcal{L}_{cal}$. We first perform a sensitivity test, in which we explore the impact of $\mathcal{L}_{cal}$ on the overall performance. Then we evaluate the incremental weighting performance of the calibration term in Appendix A.4, which investigates the effectiveness of $\mathcal{L}_{cal}$ at the early stage of fine-tuning. We also conduct an ablation study on the impact of active LoRA experts in Appendix A.5.
$\beta$ 00 $0.2$ $0.5$ $0.8$ $1$ $1.2$ $1.5$ $1.8$ $2$
ACC $\uparrow$ $75.5_{2.91}$ $76.0_{0.6}$ $75.9_{0.31}$ $76.6_{0.4}$ $77.3_{1.36}$ $\mathbf{77.9_{0.6}}$ $77.2_{0.91}$ $76.7_{0.82}$ $76.6_{0.92}$
ECE $\downarrow$ $12.9_{1.99}$ $7.47_{0.78}$ $7.82_{0.93}$ $5.94_{1.16}$ $\mathbf{5.0_{1.15}}$ $6.49_{2.65}$ $9.19_{0.86}$ $8.11_{1.43}$ $6.65_{1.85}$
Sensitivity Test on Calibration Term. To further understand the effectiveness of the calibration loss, we perform a sensitivity test of the coefficient $\beta$ in Equation 14. This evaluates how our proposed calibration of parameter mixtures affect the overall model prediction and uncertainty quantification capabilities. We evaluate $\beta$ values ranging from 00 to $2$, where $\beta=0$ resembles the original MixLoRA method.
6 Discussion & Conclusion
In this work, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which addresses the overconfidence issues commonly encountered during fine-tuning of large language models. We present a functional perspective on quantifying uncertainty in LLMs and utilize it for uncertainty-calibrated fine-tuning. By incorporating functional-level uncertainty quantification with a mixture-of-experts framework, our proposed uncertainty-calibrated training loss effectively addresses the challenge of overconfidence in fine-tuned LLMs by significantly improving uncertainty calibration while maintaining high accuracy. Our evaluations demonstrate that UQ4CT reduces the Expected Calibration Error by more than 25% without compromising accuracy across a variety of downstream tasks, including common-sense and domain-specific reasoning, under in-distribution and out-of-distribution scenarios.
Acknowledgement
Appendix A Appendix
A.1 Theoretical Derivation of the Method
In this section, we provide complete theoretical statements and proofs that are used in Sec. 3 to derive our method.
Fact A.1 (Model Perturbation Structure, Restatement of Fact 3.1).
Assume that in the residual connection architecture in each layer: $g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=\mathbf{h}^{\ell}+E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\mathbf{h}^{\ell})\big{)}$, the Lipschitzness of the residual term $E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\mathbf{h}^{\ell})\big{)}$ is much smaller as compared to that of $\mathbf{h}^{\ell}$ itself: $\left\|E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\hat{\mathbf{h}}^{\ell})\big{)}-E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\mathbf{h}^{\ell})\big{)}\right\|=o\left(\|\hat{\mathbf{h}}^{\ell}-\mathbf{h}^{\ell}\|\right)$. Under this regularity assumption, perturbations $\Delta f(x)$ to the model $f(x)$ approximately decomposes as:
$\Delta f(x)\approx\sum_{\ell=1}^{L}\sum_{k^{\ell}=1}^{K}\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell}).$
Proof of Fact 3.1 and A.1.
In each layer, we can decompose the perturbation to the output as follows:
$\Delta\mathbf{h}^{\ell+1}=\sum_{k^{\ell}=1}^{K}\left[\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})+\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot\Delta g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\right],\quad\forall\ell=1,\dots,L,$
where the input $\mathbf{h}^{1}=x$ and the output $f(x)=\mathbf{h}^{L+1}$.
Due to the residual connection architecture and our assumption on the regularity of the residual term, we have:
$\Delta g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=\Delta\mathbf{h}^{\ell}+\Delta E_{k^{\ell}}^{\ell}\left(f_{trans}^{\ell}(\mathbf{h}^{\ell})\right)=\Delta\mathbf{h}^{\ell}+o\left(\Delta\mathbf{h}^{\ell}\right).$
Hence,
$\Delta\mathbf{h}^{\ell+1}\approx\sum_{k^{\ell}=1}^{K}\left(\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})+\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot\Delta\mathbf{h}^{\ell}\right)=\sum_{k^{\ell}=1}^{K}\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})+\Delta\mathbf{h}^{\ell}.$
Expanding this recursion, the output perturbation can be approximated as:
$\Delta f(x)=\Delta\mathbf{h}^{L+1}\approx\sum_{\ell=1}^{L}\sum_{k^{\ell}=1}^{K}\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell}).$
Proposition A.2 (Calibration Loss, Restatement of Proposition 3.2).
Let the calibration risk be defined as the expectation of the calibration loss over the data distribution:
¯$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
If this calibration risk is optimized over the data distribution $\mathcal{D}$, then the optimal solution is $\mathrm{FLU}(x)=\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))$; that is, the optimally trained FLU corresponds to the probability that the model’s prediction is correct.
Proof of Proposition 3.2 and A.2.
Expanding the calibration risk, we have:
¯$\bar{\mathcal{L}}_{\mathrm{cal}}$ $=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}$
$=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left[\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}\right]^{2}$
$\quad-2\,\mathbb{E}_{x}\mathbb{E}_{y^{*}|x}\left[\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}\cdot\mathrm{FLU}(x)\right]+\mathbb{E}_{x}\left[\mathrm{FLU}(x)^{2}\right]$
$=C+\mathbb{E}_{x}\left[\mathrm{FLU}(x)^{2}-2\,\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))\cdot\mathrm{FLU}(x)\right],$
where $C$ is a constant independent of $\mathrm{FLU}(x)$. Thus, the calibration risk is minimized when $\mathrm{FLU}(x)=\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))$. ∎
A.2 Load Balancing Loss
where
A.3 Experimental Results with Mistral-7B
@@
Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintains similar accuracy.
@@
A.4 Incremental Weighting on Calibration Term
@@
As shown in Table 3, the incremental loss has significantly worse ECE performance across all tasks. This demonstrates the advantage of uncertainty calibration even in the early stage. In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high functional-level uncertainty. Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
A.5 Deciding Number of Active Experts
@@
A.6 Training Details
We train our model with total of $8$ LoRA experts, and select $2$ experts with the highest probability. For each expert, we use $rank=16$ and $alpha=32$. We use batch size of $16$ to train our model. For climate task, we set the learning rate to $45e-$ 4 and dropout rate to $0.1$ to incorporate the small dataset size. For other tasks, we use $42e-$ 4 as our learning rate with dropout $0.05$. We use AdamW as our optimizer and a cutoff length of $512$ for prompts during training. Our model is trained on A100 GPU, with 20GB GPU memory consumption per task. Training time is from $25$ to $50$ minutes depending on the task.
@@ A.6 Training Details
A.7 Expected Calibration Error
@@
In this paper, we use an ECE bin size of $15$, following the experiment setup in Laplace-LoRA [Yang et al., 2024a].
A.8 MMLU Distribution Shift Dataset Composition
@@
A.9 Prompt Perturbation Comparison
Here, we compare our method with SPUQ [Gao et al., 2024], which perturbs the prompt, aggregates predictions and confidences to measure uncertainty. We test SPUQ and UQ4CT with LLama2-7b as the base model. As shown in Table 6, SPUQ’s large ECE values suggest that simply aggregating predictions from perturbed prompts does not adequately calibrate model confidence, highlighting the limitations of prompt perturbation as an uncertainty quantification strategy for LLMs, especially for smaller models.
Table 6: Performance comparison of UQ4CT and SPUQ across five tasks. UQ4CT achieves higher accuracy and substantially lower ECE than SPUQ.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
ACC $\uparrow$ SPUQ $62.20$ $75.14$ $45.99$ $59.00$ $60.64$
UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
ECE $\downarrow$ SPUQ $30.47$ $13.13$ $24.74$ $20.00$ $19.98$